#31 3points Time Remaining 1 hour 15 minutes 12 seconds 01:15:12 Item 31 Item 31 3...

#31

3points

Time Remaining 1 hour 15 minutes 12 seconds

01:15:12

Item 31

Item 31 3 points

Time Remaining 1 hour 15 minutes 12 seconds

01:15:12

A company's flexible budget for 15,000 units of production showed total contribution margin of $76,500 and fixed costs, $37,500. The operating income expected if the company produces and sells 18,000 units is:

Multiple Choice

  • $54,300.

  • $98,700.

  • $22,200.

  • $39,000.

  • $25,735.
Answer #1

contribution margin per unit=(76500/15000)=$5.1 per unit

Hence total contribution margin at 18000 units=(18000*5.1)=$91800

Less:Fixed costs=($37500)

Operating income expected=$54300

NOTE:Total fixed costs and variable cost per unit do not change with change in units